OpenStudy (marissalovescats):

Determine whether Rolle's Theorem can be applied to f(x) on the interval [-1,1]. If Rolle's Theorem can be applied, find all values, c, in the interval such that f'(c)=0. If Rolle's Theorem cannot be applied, state why.

OpenStudy (marissalovescats):

\[f(x) = \frac{ x^3+x }{ x}\]

OpenStudy (marissalovescats):

And I know Rolle's Theorem: f(a) = f(b) then f'(c)=0 Idk why I got this problem wrong but I did

OpenStudy (lyrae):

"If a real-valued function f is continuous on a closed interval [a, b], differentiable on the open interval (a, b), and f(a) = f(b) ..." -Wikipedia f(x) is't continous on the intreval [-1,1] see x=0.

OpenStudy (marissalovescats):

Okay well all of that is true